OpenStudy (sidmanfu):

A father has a son and daughter. The father is five times as old as the daughter and the son is 2 years older than the daughter. In 10 years, the sum of their ages will be 81. How old are they now? Solve this problem by setting up an equation.

OpenStudy (anonymous):

write that as an equation , and try to solve it

OpenStudy (sidmanfu):

im not really sure how to make it into an equation.

OpenStudy (anonymous):

lets name father ( f ) , daughter ( d ) and the son ( s ) f = 5 d s = d + 2 (f+10) + ( s + 10 ) + (d + 10 ) = 81 try to solve it !

OpenStudy (anonymous):

can u solve it ?

OpenStudy (sidmanfu):

Im stuck lol

OpenStudy (anonymous):

so, now that you got the equation, to solve it you need to keep only one unknown number f = 5d s = d + 2 , any idea ?

OpenStudy (sidmanfu):

sorry but I have no idea

OpenStudy (anonymous):

how much is d in terms of s ?

OpenStudy (sidmanfu):

added by 2...

OpenStudy (anonymous):

thats s in terms of d , we need d in terms of s ))

OpenStudy (sidmanfu):

ohh is it 2 year younger?

OpenStudy (anonymous):

yes so , d = s - 2 , so f = 5 ( s - 2 ) right ?

OpenStudy (sidmanfu):

yes thats right

OpenStudy (anonymous):

now simplify that f = 5( s - 2 ) , and get me s in the terms of f

OpenStudy (sidmanfu):

so is \[f = 5s - 10\] ?

OpenStudy (anonymous):

yes , now get s in terms of f s = ??